Third National Bank has reserves of $10,000 and checkable deposits of $100,000. The reserve ratio is 10 percent. Households deposit $15,000 in currency into the bank and that currency is added to reserves.

What level of excess reserves does the bank now have?
Answer #1

ANSWER:

Checkable deposits = 100,000 + 15,000 = 115,000

Required reserves = reserve ratio * chheckable deposits = 10% * 115,000 = 11,500

Excess reserves = actual reserves - required reserves = (10,000 + 15,000) - 11,500 = 25,000 - 11,500 = 13,500

Excess reserves = $13,500